Operations Report – April 24th

Crew 193 Operations Report 24Apr2018

SOL: 2

Name of person filing report: David Attig

Non-nominal systems: One piece EVA suit #10 and two piece EVA suit # 6 have issues

Notes on non-nominal systems: One two piece suit (6) has straps broken holding on neck ring, Visor on suit 10 is heavily scratched and deemed unuseable. See Sol Summary for more details.

Generator (hours run): 12

Solar— SOC 86 % (Before generator is run at night)

Diesel – 75 %

Propane – 60 psi.

ATV Fuel (Ethanol free gasoline) – 0.75 gallons.

Water (trailer) – 0 gallons.

Water (storage) – 550 gallons

Water (static) – 410 gallons

Trailer to Static Pump used – no

Water (loft) – Static to Loft Pump used – yes

Water Meter: 134008.9

Toilet tank emptied: yes
Date last emptied: April 24

ATV’s (Honda, 300, 350.1, 350.2, 350.3):
Honda Used?: Yes
300 Used? Yes
350.1 Used? No
350.2 Used? Yes
350.3 Used? No

ATV Oil Checked?: Yes

ATV Oil Added?: No

ATV Fuel Used: less than 1 Gallon

# Hours the ATVs were Used today: 0.5

Notes on ATVs: With 350.2, throttle cable gets pulled when handlebar is moved all the way to the left. Suggest not using this ATV for future (provided sufficient fuel in others).

Deimos rover used: Yes
Hours: 124.0
Beginning charge: 99%
Ending charge: 98 %
Currently charging: Yes

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: Unknown
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Spirit rover used: No
Hours: 33.7
Beginning charge: 100 %
Ending charge: 100 %
Currently charging: No

Opportunity rover used: Out Of Service
Hours: 31
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Curiosity rover used: No
Hours: 23
Beginning charge: 100 %
Ending charge: 100 %
Currently charging: No

Hab Car used?: No
Where was Hab Car used?: N/A
Why was Hab Car used?: N/A

General notes and comments:

Summary of internet: By 8AM again today, we have used half of our allotted daily internet. Is there a way to monitor our internet over time without constantly logging on to check it? That way we could monitor usage and know if there is a peak time that it is being used.

Summary of suits and radios: Older style (silver front) radios appear to not transmit well enough to use in EVA. One two piece suit broke during EVA today, and 1 one piece suit was deemed unuseable due to scratches in the visor.

Summary of Hab operations: See Sol Summary report.

Summary of GreenHab operations: See GreenHab report

Summary of ScienceDome operations: See Science report.

Summary of RAM operations: We cut a board to use for a new sign in front of MDRS.

Summary of health and safety issues: None to report.

Questions, concerns and requests to Mission Support: See Sol Summary.

Copyright © The Mars Society. All rights reserved. | Main Site